Required information

[The following information applies to the questions displayed below.]

Smith-Kline Company maintains inventory records at selling prices as well as at cost. For 2021, the records indicate the following data:

($ in 000s)
Cost Retail
Beginning inventory $ 85 $ 130
Purchases 676 1,011
Freight-in on purchases 35
Purchase returns 1 2
Net markups 3
Net markdowns 7
Net sales 921

Required:
1. Use the retail method to approximate cost of ending inventory valued under Average cost method. (Enter your answer in thousands including 2 decimal places, i.e. 12,550 would be 12.55.)

2. Use the retail method to approximate cost of ending inventory valued under Conventional method. (Enter your answer in thousands including 2 decimal places, i.e. 12,550 would be 12.55)
Answer #1
Cost of ending inventory :
($ 000s)
1. Average cost ( use cost to retail ratio B above) = 214 * 70.0441% =     149.89
2. Conventional ( use cost to retail ratio A above) = 214 * 69.1471 % =     147.97

Workings:

($ 000s) ($ 000s)
Cost Retail
Beginning inventory       85.00     130.00
Purchases     676.00 1,011.00
Less : Purchase return        -1.00        -2.00
Freight in on purchases       35.00
Net markups         3.00
Merchandise available for sale     795.00 1,142.00
Cost to retail ratio = Cost / Retail = 795 / 1142 = 69.1471 % (A)
Less :
Net markdowns        -7.00
Merchandise available for sale     795.00 1,135.00
Cost to retail ratio = Cost / Retail = 795 / 1135 = 70.0441 % (B)
Less : Net Sales    -921.00
Ending inventory at retail     214.00